# Three-body forces and Efimov physics in nuclei and atoms

### v1

@@ 1 Introduction
All visible matter in the universe is organized into resolution-dependent hierarchical structures. A precise description of the interactions at each hierarchical level starting from elementary quarks to composite systems like hadrons, nuclei, atoms and molecules may help to improve our understanding of the structure and dynamics of strongly interacting matter. Generally, effective interactions at all hierarchical levels are dominated by pairwise forces acting between two constituent particles. Meanwhile, recent advances in computational, theoretical and experimental techniques allow one to go beyond the two-body-force level and to quantitatively probe three-body interactions, which are known to figure importantly in atomic and nuclear systems Hammer:2012id .
@@ 2.1 Three-nucleon force and its importance in nuclear phenomena
Since Yukawa’s meson theory proposed in 1935 Yukawa193548 , the nuclear force has been modeled in terms of meson exchange interactions between nucleons. Beside the dominant two-nucleon forces, the three nucleon forces ($3N$Fs) have attracted an increasing attention in the last two decades. The development of high-precision two-nucleon potentials in the 90s of the last century, coupled with advances in ab initio few-body calculations based on these interactions, have confirmed the important role played by $3N$Fs in various nuclear phenomena. Following the seminal work by Fujita and Miyazawa, a number of the $3N$F models utilizing the longest-range $2\pi$-exchange mechanism have been developed such as, e.g., the Tucson-Melbourne 99 tm99 , the Urbana IX PhysRevC.56.1720 $3N$F models. A new impetus to study $3N$Fs has come from chiral perturbation theory, the low-energy effective field theory of QCD Weinberg:1990rz ; kolk1994 ; Epelbaum:2002vt .
@@ 2.2.2 Chiral EFT for nuclear systems
Clearly, the singular short-distance nature of $V_{NN}^{1\pi}(\vec{q}\,)$ is unphysical and represents an artifact of using the low-momentum approximation in Eq. (2) beyond its validity range (i.e., at short distances or large values of the momentum transfer). Renormalization of the Schrödinger equation in EFTs with singular interactions like $V_{NN}^{1\pi}(\vec{q}\,)$ can be achieved in the way compatible with the weak-interaction limit by introducing a finite cutoff $\Lambda$ of the order of the pertinent hard scale, $\Lambda\sim\Lambda_{b}$ Lepage:1997cs . At each order of ChEFT and for every value of $\Lambda$, renormalization is carried out implicitly by expressing bare LECs from $\mathcal{L}_{NN}$, $\mathcal{L}_{NNN}$, $\ldots$, in terms of measurable quantities. In practice, this is achieved by tuning few-nucleon contact interactions to experimental data. The residual dependence of the renormalized results on the cutoff $\Lambda$ serves as a measure of the neglected contributions of terms beyond the EFT truncation level. It is expected to decrease with the chiral order and provides an a posteriori consistency check. In Refs. Gasparyan:2021edy ; Gasparyan:2023rtj , the finite-$\Lambda$ formulation of ChEFT was formally proven to be renormalizable (in the EFT sense) up to next-to-leading order $Q^{2}$ (NLO). Here and in what follows, we restrict ourselves to the finite-cutoff formulation of ChEFT as described in detail in Ref. Epelbaum:2019kcf . A pedagogical introduction into the considered framework can be found in Sec. 6.3 of Ref. Gross2022hyw , while a discretized (lattice) formulation is described in Ref. Lahde:2019npb .
@@ 2.2.3 Current status of the $NN$ potentials
Presently, the $NN$ force has been worked out completely up through fifth order $Q^{5}$ (i.e., N4LO) including isospin-breaking corrections due to $m_{u}\neq m_{d}$ and QED effects. It involves the one- two- and three-pion exchange potentials, supplemented by short-range NN interactions from $\mathcal{L}_{NN}^{(0)}$, $\mathcal{L}_{NN}^{(2)}$, and $\mathcal{L}_{NN}^{(4)}$, which depend on $2$, $7$ and 15151515111The above numbers refer to isospin-invariant contact interactions. LECs, respectively, that need to be tuned to $NN$ data. Notice that $3$ out of $15$ contact interactions at N4LO contribute only to the off-shell part of the $NN$ potential and thus cannot be determined from two-nucleon scattering data.
@@
where the ellipses denote subtractions in the form of $NN$ contact interactions, chosen in such a way that the corresponding $r$-space potentials vanish at the origin. In the second equation, we have used a dispersive representation of the TPEP Kaiser:1997mw , where the spectral function $\rho_{2\pi}(\mu)$ is obtained from an analytic continuation of the unregularized potential $V_{NN}^{2\pi}(q)$ to imaginary values of $q$ via $\rho_{2\pi}(\mu):=\Im\big{[}V_{NN}^{2\pi}(0^{+}-i\mu)\big{]}$. Equations (3) and (4) show that all finite-$\Lambda$ artifacts stemming from the $1/\Lambda$-expansion of the regularized OPEP and TPEP have the form of short-range contact interactions. The latter are regularized in Ref. Reinert:2017usi using a Gaussian non-local regulator $\exp[-(p^{2}+p^{\prime 2})/\Lambda^{2}]$. The resulting semi-local momentum-space regularized (SMS) $NN$ potentials are available at orders LO through N4LO and cutoff values of $\Lambda=400$, $450$, $500$ and $550$ MeV. The N4LO+ potentials lead to a statistically perfect description of the Granada-2013 database of mutually consistent neutron-proton and proton-proton scattering data below pion production threshold NavarroPerez:2013usk and show, as expected, a very weak residual $\Lambda$-dependence. In Ref. Reinert:2020mcu , the SMS $NN$ potential of Ref. Reinert:2017usi was updated to include all relevant isospin-breaking contributions up through N4LO, and it was used for a precision determination of the $\pi N$ coupling constants from $NN$ scattering and to perform a full fledged partial wave analysis of $NN$ data (including a selection of mutually compatible data), see Ref. Epelbaum:2022cyo for details and Fig. 2 for representative examples
@@ 2.2.4 Chiral expansion of the $3N$F
So far we have left open the question of regularization of the $3N$F. As already pointed out in Sec. 2.2.2, ChEFT allows one to derive low-momentum approximations of the nuclear forces, but the method looses its validity for large momenta $|\vec{p}\,|\gtrsim\Lambda_{b}$. For example, the ChEFT predictions for the function ${\cal{A}}(q_{2})$ in Eqs. (2.2.4), (9) and (2.2.4) behave at large $q_{2}$-values as ${\cal{A}}_{[Q^{3}]}(q_{2})\sim q_{2}^{2}$, ${\cal{A}}_{[Q^{4}]}(q_{2})\sim q_{2}^{3}$ and ${\cal{A}}_{[Q^{5}]}(q_{2})\sim q_{2}^{4}\,\log(q_{2})$, respectively. While the expansion for ${\cal{A}}(q_{2})$ appears to converge well within the applicability range of ChEFT, with the order-$Q^{5}$ result providing a tiny correction as shown in the left panel of Fig. 4, the N4LO contribution starts exceeding the N3LO correction at $q_{2}\sim 780$ MeV, signalling the breakdown of the expansion at large momenta. As explained in Sec. 2.2.2, the expressions for nuclear forces derived in ChEFT have to be regularized by introducing a finite cutoff $\Lambda\sim\Lambda_{b}$ to remove their uncontrolled behavior at large momenta and make the $A$-body Schrödinger equation well defined. In the $NN$ sector, this can be achieved, e.g., by simply multiplying the unregularized potentials with some regulator functions. Unfortunately, following this same strategy for the $3N$F was found in Refs. Epelbaum:2019kcf ; Epelbaum:2019jbv ; Epelbaum:2022cyo to violate the chiral symmetry. The problem can be best illustrated with the $3N$F of type (b) stemming from the first of the two diagrams shown in the second line and third column of Fig. 3. When calculating the scattering amplitude, the irreducible $3N$F is supplemented with the corresponding reducible contributions stemming from the iteration of the Faddeev equation. This reducible contribution involves a linearly divergent piece $\propto\Lambda$, whose structure violates the chiral symmetry. Accordingly, the amplitude cannot be renormalized by re-adjusting the LEC $c_{D}$, and the whole ChEFT expansion for $3N$ scattering breaks down. As explained in Refs. Epelbaum:2019kcf ; Epelbaum:2019jbv ; Epelbaum:2022cyo , the problem is caused by mixing of two different regularization schemes, namely DimReg when deriving the $3N$F and cutoff regularization in the Schrödinger equation. The same issue affects loop contributions to the nuclear current operators Krebs:2019ddp ; Krebs:2020pii . This shows that a consistent regularization of the $3N$F and nuclear currents cannot be achieved by simply multiplying the corresponding potentials derived using DimReg with some cutoff functions. Rather, $3N$Fs, 4NFs and nuclear currents starting from N3LO need to be re-derived using a cutoff instead of DimReg.
@@ 2.2.5 Selected applications of the lading $3N$F
In a series of papers by the LENPIC Collaboration, the semi-local coordinate-space regularized (SCS) $NN$ potentials of Refs. Epelbaum:2014efa ; Epelbaum:2014sza at all available orders from LO to N4LO were employed to calculate $Nd$ scattering observables and properties of selected nuclei up to 48Ca LENPIC:2015qsz ; Maris:2016wrd ; LENPIC:2018lzt . While these calculations did not include the $3N$F and thus should be regarded as incomplete starting from N2LO, they have yielded a number of interesting observations. In particular, calculations beyond the last complete order NLO were found to differ from experimental data well outside the estimated truncation uncertainties, thus providing unambiguous evidence for missing $3N$Fs. Moreover, the amount of deviations between theory and experimental data was found to be consistent with estimations based on the power counting.
@@
Figure 5: ChEFT predictions for the differential cross section, deuteron vector analyzing power $A_{y}^{d}$ and deuteron tensor analyzing powers $A_{xz}$ and $A_{xx}$ in elastic neutron-deuteron scattering at $E_{\rm lab}=135$ MeV. Dark-shaded orange and green bands show the NLO and N2LO results at the $1\sigma$ confidence level, respectively, while the corresponding light-shaded bands show the $2\sigma$-intervals. Experimental data are pd elastic scattering data from Ref. sekiguchi2002 . Dashed lines in the middle of green bands are the actual N2LO predictions. Dotted lines are obtained using the NN interaction at the highest available order N4LO+, supplemented with the N2LO $3N$F (with the appropriately re-adjusted LECs $c_{D}$ and $c_{E}$). In all calculations, the cutoff is chosen to be $\Lambda=450$ MeV.
In Fig. 5, we show selected results for $Nd$ elastic scattering observables at $E_{\rm lab}=135$ MeV, which may serve as representative examples. Given that the LECs $c_{D}$ and $c_{E}$ are fixed from the 3H binding energy and the differential cross section minimum at $E_{\rm lab}=70$ MeV, the shown results are to be regarded as predictions. The experimental data from Ref. sekiguchi2002 are mostly in agreement with the calculations (within errors), but the N2LO truncation uncertainty at this moderate energy appears to be rather large. The description of $Nd$ data at N2LO is qualitatively similar to the one for proton-proton scattering as a comparable energy, shown in Fig. 2. Based on the results in the $NN$ system, it is expected that taking into account the $3N$F up through N4LO would allow one to achieve a precise description of $Nd$ scattering data, comparable to that of the neutron-proton and proton-proton data reported in Refs. Reinert:2017usi ; Reinert:2020mcu .
@@ 2.3 Experimental studies of three-nucleon forces
The importance of $3N$Fs in the continuum spectrum was shown, for the first time, in nucleon–deuteron ($Nd$) elastic scattering at the end of the 1990s wit98 . $3N$Fs were found to lead to pronounced effects around the cross-section minimum occurring at the values of the c.m. scattering angle of $\theta_{c.m.}\approx 120^{\circ}$ for incident energies above $60~{}\rm MeV/nucleon$. Since then, proton-deuteron($pd$) and neutron-deuteron($nd$) scattering experiments at 60–300 MeV/nucleon have been performed at the facilities, e.g. RIKEN, RCNP, KVI, IUCF, TSL, and LANSCE, providing precise data of the cross sections as well as various types of spin observables Kalantar-Nayestanaki_2012 .
@@ 3.3 Artificial control of three-body forces
Three-body forces not only naturally appear in low dimensions, but they can also be controlled artificially with cold atoms. Accordingly, it is even possible to make three-body forces dominate over two-body forces. While several such schemes have been proposed theoretically PhysRevA.89.053619 ; PhysRevLett.112.103201 ; PhysRevA.90.021601 ; PhysRevA.93.043616 , we here introduce the simple and versatile one proposed in Ref. PhysRevA.90.021601 , which employs two hyperfine spin components of bosons in an optical lattice. When the two components are coupled by a nearly resonant field, the system in the tight-binding approximation is described by

### v2

@@ 1 Introduction
All visible matter in the universe is organized into resolution-dependent hierarchical structures. A precise description of the interactions at each hierarchical level starting from elementary quarks to composite systems like hadrons, nuclei, atoms and molecules may help to improve our understanding of the structure and dynamics of strongly interacting matter. Generally, effective interactions at all hierarchical levels are dominated by pairwise forces acting between two constituent particles. Meanwhile, recent advances in computational, theoretical and experimental techniques allow one to go beyond the two-body-force level and to quantitatively probe three-body interactions, which are known to figure importantly in atomic and nuclear systems gibson1988three ; Hammer:2012id .
@@ 2.1 Three-nucleon force and its importance in nuclear phenomena
Since Yukawa’s meson theory proposed in 1935 Yukawa193548 , the nuclear force has been modeled in terms of meson exchange interactions between nucleons. Beside the dominant two-nucleon forces, the three nucleon forces ($3N$Fs) have attracted an increasing attention in the last three decades. The development of high-precision two-nucleon potentials in the 90s of the last century, coupled with advances in ab initio few-body calculations based on these interactions, have confirmed the important role played by $3N$Fs in various nuclear phenomena. Following the seminal work by Fujita and Miyazawa, a number of the $3N$F models utilizing the longest-range $2\pi$-exchange mechanism have been developed such as, e.g., the Tucson-Melbourne 99 tm99 , the Urbana IX PhysRevC.56.1720 $3N$F models. A new impetus to study $3N$Fs has come from chiral perturbation theory, the low-energy effective field theory of QCD Weinberg:1990rz ; kolk1994 ; Epelbaum:2002vt .
@@ 2.2.2 Chiral EFT for nuclear systems
Clearly, the singular short-distance nature of $V_{NN}^{1\pi}(\vec{q}\,)$ is unphysical and represents an artifact of using the low-momentum approximation in Eq. (2) beyond its validity range (i.e., at short distances or large values of the momentum transfer). Renormalization of the Schrödinger equation in EFTs with singular interactions like $V_{NN}^{1\pi}(\vec{q}\,)$ can be achieved in the way compatible with the weak-interaction limit by introducing a finite cutoff $\Lambda$ of the order of the pertinent hard scale, $\Lambda\sim\Lambda_{b}$ Lepage:1997cs . At each order of ChEFT and for every value of $\Lambda$, renormalization is carried out implicitly by expressing bare LECs from $\mathcal{L}_{NN}$, $\mathcal{L}_{NNN}$, $\ldots$, in terms of measurable quantities. In practice, this is achieved by tuning few-nucleon contact interactions to experimental data. The residual dependence of the renormalized results on the cutoff $\Lambda$ serves as a measure of the neglected contributions of terms beyond the EFT truncation level. It is expected to decrease with the chiral order and provides an a posteriori consistency check. In Refs. Gasparyan:2021edy ; Gasparyan:2023rtj , the finite-$\Lambda$ formulation of ChEFT was formally proven to be renormalizable (in the EFT sense) up to next-to-leading order $Q^{2}$ (NLO). Here and in what follows, we restrict ourselves to the finite-cutoff formulation of ChEFT as described in detail in Ref. Epelbaum:2019kcf . A pedagogical introduction into the considered framework can be found in Sec. 6.3 of Ref. Gross2022hyw , while a discretized (lattice) formulation is described in Ref. Lahde:2019npb . For a collection of different views on renormalization and power counting issues in nuclear effective field theories see Ref. Tews:2022yfb and references therein.
@@ 2.2.3 Current status of the $NN$ potentials
Presently, the $NN$ force has been worked out completely up through fifth order $Q^{5}$ (i.e., N4LO) including isospin-breaking corrections due to $m_{u}\neq m_{d}$ and QED effects. It involves the one- two- and three-pion exchange potentials, supplemented by short-range NN interactions from $\mathcal{L}_{NN}^{(0)}$, $\mathcal{L}_{NN}^{(2)}$, and $\mathcal{L}_{NN}^{(4)}$, which depend on $2$, $7$ and 15151515111The above numbers refer to isospin-invariant contact interactions. LECs, respectively, that need to be tuned to $NN$ data. Notice that $3$ out of $15$ contact interactions at N4LO contribute only to the off-shell part of the $NN$ potential and thus cannot be determined from two-nucleon scattering data. In the NN potentials of Refs. Reinert:2017usi ; Reinert:2020mcu , these off-shell LECs are set to zero. To analyze $A\geq 3$ nuclear systems at N3LO, these LECs will have to be determined from experimental data beyond the NN system, such as, e.g., three-nucleon scattering data.
@@
where $\ldots$” denote subtractions in the form of $NN$ contact interactions, chosen in such a way that the corresponding $r$-space potentials vanish at the origin. In the second equation, we have used a dispersive representation of the TPEP Kaiser:1997mw , where the spectral function $\rho_{2\pi}(\mu)$ is obtained from an analytic continuation of the unregularized potential $V_{NN}^{2\pi}(q)$ to imaginary values of $q$ via $\rho_{2\pi}(\mu):=\Im\big{[}V_{NN}^{2\pi}(0^{+}-i\mu)\big{]}$. Equations (3) and (4) show that all finite-$\Lambda$ artifacts stemming from the $1/\Lambda$-expansion of the regularized OPEP and TPEP have the form of short-range contact interactions. The latter are regularized in Ref. Reinert:2017usi using a Gaussian non-local regulator $\exp[-(p^{2}+p^{\prime 2})/\Lambda^{2}]$. The resulting semi-local momentum-space regularized (SMS) $NN$ potentials are available at orders LO through N4LO and cutoff values of $\Lambda=400$, $450$, $500$ and $550$ MeV. The N4LO+ potentials lead to a statistically perfect description of the Granada-2013 database of mutually consistent neutron-proton and proton-proton scattering data below pion production threshold NavarroPerez:2013usk and show, as expected, a very weak residual $\Lambda$-dependence. In Ref. Reinert:2020mcu , the SMS $NN$ potential of Ref. Reinert:2017usi was updated to include all relevant isospin-breaking contributions up through N4LO, and it was used for a precision determination of the $\pi N$ coupling constants from $NN$ scattering and to perform a full fledged partial wave analysis of $NN$ data (including a selection of mutually compatible data), see Ref. Epelbaum:2022cyo for details and Fig. 2 for representative examples
@@ 2.2.4 Chiral expansion of the $3N$F
So far we have left open the question of regularization of the $3N$F. As already pointed out in Sec. 2.2.2, ChEFT allows one to derive low-momentum approximations of the nuclear forces, but the method loses its validity for large momenta $|\vec{p}\,|\gtrsim\Lambda_{b}$. For example, the ChEFT predictions for the function ${\cal{A}}(q_{2})$ in Eqs. (2.2.4), (9) and (2.2.4) behave at large $q_{2}$-values as ${\cal{A}}_{[Q^{3}]}(q_{2})\sim q_{2}^{2}$, ${\cal{A}}_{[Q^{4}]}(q_{2})\sim q_{2}^{3}$ and ${\cal{A}}_{[Q^{5}]}(q_{2})\sim q_{2}^{4}\,\log(q_{2})$, respectively. While there is an indication that the expansion for ${\cal{A}}(q_{2})$ converges well within the applicability range of ChEFT, with the order-$Q^{5}$ result providing a tiny correction as shown in the left panel of Fig. 4, the N4LO contribution starts exceeding the N3LO correction at $q_{2}\sim 780$ MeV, signalling the breakdown of the expansion at large momenta. As explained in Sec. 2.2.2, the expressions for nuclear forces derived in ChEFT have to be regularized by introducing a finite cutoff $\Lambda\sim\Lambda_{b}$ to remove their uncontrolled behavior at large momenta and make the $A$-body Schrödinger equation well defined. In the $NN$ sector, this can be achieved, e.g., by simply multiplying the unregularized potentials with some regulator functions. Unfortunately, following this same strategy for the $3N$F was found in Refs. Epelbaum:2019kcf ; Epelbaum:2019jbv ; Epelbaum:2022cyo to violate the chiral symmetry. The problem can be best illustrated with the $3N$F of type (b) stemming from the first of the two diagrams shown in the second line and third column of Fig. 3. When calculating the scattering amplitude, the irreducible $3N$F is supplemented with the corresponding reducible contributions stemming from the iteration of the Faddeev equation. This reducible contribution involves a linearly divergent piece $\propto\Lambda$, whose structure violates the chiral symmetry. Accordingly, the amplitude cannot be renormalized by re-adjusting the LEC $c_{D}$, and the whole ChEFT expansion for $3N$ scattering breaks down. As explained in Refs. Epelbaum:2019kcf ; Epelbaum:2019jbv ; Epelbaum:2022cyo , the problem is caused by mixing of two different regularization schemes, namely DimReg when deriving the $3N$F and cutoff regularization in the Schrödinger equation. The same issue affects loop contributions to the nuclear current operators Krebs:2019ddp ; Krebs:2020pii . This shows that a consistent regularization of the $3N$F and nuclear currents cannot be achieved by simply multiplying the corresponding potentials derived using DimReg with some cutoff functions. Rather, $3N$Fs, 4NFs and nuclear currents starting from N3LO need to be re-derived using a cutoff instead of DimReg.
@@ 2.2.5 Selected applications of the lading $3N$F
In a series of papers by the LENPIC Collaboration, the semi-local coordinate-space regularized (SCS) $NN$ potentials of Refs. Epelbaum:2014efa ; Epelbaum:2014sza at all available orders from LO to N4LO were employed to calculate $Nd$ scattering observables and properties of selected nuclei up to 48Ca LENPIC:2015qsz ; Maris:2016wrd ; LENPIC:2018lzt . While these calculations did not include the $3N$F and thus should be regarded as incomplete starting from N2LO, they have yielded a number of interesting observations. In particular, calculations beyond the last complete order NLO were found to differ from experimental data well outside the estimated truncation uncertainties at the corresponding expansion orders, thus providing unambiguous evidence for missing $3N$Fs. Moreover, the amount of deviations between theory and experimental data was found to be consistent with estimations based on the power counting.
@@
Figure 5: ChEFT predictions for the differential cross section, deuteron vector analyzing power $A_{y}^{d}$ and deuteron tensor analyzing powers $A_{xz}$ and $A_{xx}$ in elastic neutron-deuteron scattering at $E_{\rm lab}=135$ MeV. Dark-shaded orange and green bands show the NLO and N2LO results at the $1\sigma$ confidence level, respectively, while the corresponding light-shaded bands show the $2\sigma$-intervals. The data points are experimental $pd$ elastic scattering data taken at RIKEN and RCNP (grey circles) sekiguchi2002 ; nsakamot96 ; PhysRevLett.95.162301 , KVI (brown diamonds) PhysRevC.78.014006 ; PhysRevLett.86.5862 , and IUCF (blue triangles) PhysRevC.74.064003 . Dashed lines in the middle of green bands are the actual N2LO predictions. Dotted lines are obtained using the NN interaction at the highest available order N4LO+, supplemented with the N2LO $3N$F (with the appropriately re-adjusted LECs $c_{D}$ and $c_{E}$). In all calculations, the cutoff is chosen to be $\Lambda=450$ MeV.
In Fig. 5, we show selected results for $Nd$ elastic scattering observables at $E_{\rm lab}=135$ MeV, which may serve as representative examples. Given that the LECs $c_{D}$ and $c_{E}$ are fixed from the 3H binding energy and the differential cross section minimum at $E_{\rm lab}=70$ MeV, the shown results are to be regarded as predictions. The experimental data from Refs. sekiguchi2002 ; nsakamot96 ; PhysRevLett.95.162301 ; PhysRevC.78.014006 ; PhysRevLett.86.5862 ; PhysRevC.74.064003 are mostly in agreement with the calculations (within errors), but the N2LO truncation uncertainty at this moderate energy appears to be rather large. The description of $Nd$ data at N2LO is qualitatively similar to the one for proton-proton scattering as a comparable energy, shown in Fig. 2. Based on the results in the $NN$ system, it is expected that taking into account the $3N$F up through N4LO would allow one to achieve a precise description of $Nd$ scattering data, comparable to that of the neutron-proton and proton-proton data reported in Refs. Reinert:2017usi ; Reinert:2020mcu .
@@ 2.3 Experimental studies of three-nucleon forces
[…]
@@ 3.3 Artificial control of three-body forces
Three-body forces not only naturally appear in low dimensions, but they can also be controlled artificially with cold atoms. Accordingly, it is even possible to make three-body forces dominate over two-body forces, requiring to go beyond perturbations. While several such schemes have been proposed theoretically PhysRevA.89.053619 ; PhysRevLett.112.103201 ; PhysRevA.90.021601 ; PhysRevA.93.043616 , we here introduce the simple and versatile one proposed in Ref. PhysRevA.90.021601 , which employs two hyperfine spin components of bosons in an optical lattice. When the two components are coupled by a nearly resonant field, the system in the tight-binding approximation is described by